# Holographic Intelligence Surface Assisted Integrated Sensing and Communication

### v1

@@ II-B Communications and Radar Performances
$+\underbrace{\sum_{i=1,i\neq k}^{K+M}j\kappa Z_{0}\int_{D}G_{k}(p)\rho_{i}(p)x_{i}\mathrm{d}p}_{\rm Interference}+v_{c},~{}k=1,...,K,$
where $v_{c}$ is the additive white Gaussian noise (AWGN) with variance $\sigma_{c}^{2}$, $\kappa$ is the wavenumber, and $Z_{0}$ is the wave impedance in the free space. The communications SINR of the $k$-th user is given by
$\centering\eta_{c}(\boldsymbol{\rho};k)=\frac{\mathbf{E}\left(\left|j\kappa Z_{0}\int_{D}G_{k}(p)\rho_{k}(p)c_{k}\mathrm{d}p\right|^{2}\right)}{\mathbf{E}\left(\left|\sum_{i=1,i\neq k}^{K+M}j\kappa Z_{0}\int_{D}G_{k}(p)\rho_{i}(p)x_{i}\mathrm{d}p\right|^{2}\right)+\sigma_{c}^{2}},$ (3)
$k=1,...,K.$
In our considered HISAC system, both communication data streams and radar streams are utilized to support multi-target sensing since they are perfectly known by the $\mathcal{P}_{R}$. Considering both transmit and receive patterns, the sensing SINR is formulated in the following.
Let $G_{m}(p)\triangleq G\left(p^{\prime}_{m},p\right)$ and $G_{m}^{*}(p)\triangleq G\left(p_{m},p^{\prime}\right)$ denote the Green’s function from source $p$ to $m$-th targets and from the $m$-th targets to source $p$, respectively. The $m$-th target’s echo at $p^{\prime}$ is expressed as
$\centering y_{m}(p^{\prime})=\alpha_{m}G_{m}^{*}(p)e_{m},~{}m=1,...,M,~{}p^{\prime}\in\mathcal{P}_{R},$ (4)
@@
$\centering y_{l}=$ $\sum_{m=1}^{M}\int_{D}\tau_{l}\left(p^{\prime}\right)y_{m}\left(p^{\prime}\right)\mathrm{d}p^{\prime}+\int_{D}\tau_{l}v_{r}\left(p^{\prime}\right)\mathrm{d}p^{\prime}$ (5)
$=$ $e_{l}\int_{D}\tau_{l}\left(p^{\prime}\right)G_{l}^{*}(p)\mathrm{d}p^{\prime}+\underbrace{\sum_{m=1,m\neq l}^{M}e_{m}\int_{D}\tau_{l}\left(p^{\prime}\right)G_{m}^{*}(p)\mathrm{d}p^{\prime}}_{\rm Interference}$
$+\int_{D}\tau_{l}v_{r}\left(p^{\prime}\right)\mathrm{d}p^{\prime},~{}l=1,...,M,$
where ${v}_{r}(p^{\prime}),~{}p^{\prime}\in D$ is the AWGN at $\mathcal{P}_{R}$ with variance $\sigma_{r}^{2}$. Therefore, the corresponding sensing SINR is given by
$\centering\eta_{r}(\boldsymbol{\rho},\boldsymbol{\tau};l)=\frac{\mathbf{E}\left(\left|e_{l}\int_{D}\tau_{l}\left(p^{\prime}\right)G_{l}^{*}(p)\mathrm{d}p^{\prime}\right|^{2}\right)}{\mathbf{E}\left(\left|\sum_{m=1,m\neq l}^{M}e_{m}\int_{D}\tau_{l}\left(p^{\prime}\right)G_{m}^{*}(p)\mathrm{d}p^{\prime}\right|^{2}\right)+\sigma_{r}^{2}},$ (6)
$l=1,...,M.$
@@ II-C Problem Formulation in Continuous Domain
$\ \int_{D}\|\boldsymbol{\rho}(p)\|_{2}^{2}\mathrm{d}p\leq P_{T},$ (7c)
$\ \int_{D}\|{\tau}_{l}(p^{\prime})\|_{2}^{2}\mathrm{d}p^{\prime}=1,~{}l=1,...,M.$ (7d)
The primary goal of (7) is to balance the sensing SINR among radar targets, and (7b) is the communication constraint for multiple users, where $\Gamma_{c}$ represents the minimal required communication SINR. Concerning power constraints, (7c) and (7d) represent the power budget for the transmit pattern at $\mathcal{P}_{S}$ and the unit-power constraint for the receive pattern at $\mathcal{P}_{R}$, respectively.
Due to the infinite dimension integrals in the objective and constraints of the problem (7), classical beamforming techniques for discrete-array-based ISAC systems are not applicable. To address these infinite dimension functions in problem (7), we develop a continuous-discrete transformation inspired by the Fourier transform, detailed in the following section.
@@ III Continuous-Discrete Transformation
In this section, we describe the continuous-discrete transformation to solve the challenging problem in (7). The key idea is to use the Fourier transform to convert the pattern from the continuous domain into a discrete wavenumber domain. Specifically, we formulate a comprehensive continuous-discrete transformation for $\boldsymbol{\rho}$, which is then applied to derive the expressions for communication and sensing SINRs. Based on this transformation, we approximate the continuous pattern design in (7) into a joint transmit-receive beamforming problem in the discrete domain.
@@ III-A Comprehensive Continuous-Discrete Transformation
To overcome the integral functions in the pattern design, we provide a closed-form expression for $f_{2D}^{n}\left(G_{p^{\prime}}\right)$ in the following proposition.
Proposition 1
Let $n=(n_{x},n_{y})$ denote the order along the $x$-axis and $y$-axis, and $(s_{x},s_{y})$ be a source point at the surface $D$. Given the spatial Fourier transform function [19],
$\centering\Psi_{n}(p)=\frac{1}{\sqrt{A_{T}}}e^{-j2\pi\left(\frac{n_{x}}{L_{x}}\left(s_{x}-\frac{L_{x}}{2}\right)\right)}e^{-j2\pi\left(\frac{n_{y}}{L_{y}}\left(s_{y}-\frac{L_{y}}{2}\right)\right)},$ (10)
a closed-form expression for $f_{2D}^{n}\left(G_{p^{\prime}}\right)$ of the target $(r,\theta,\psi)$ on the far-field is
$\centering f_{2D}^{n}\left(G_{p^{\prime}}\right)=\frac{e^{j\kappa r}\sqrt{A_{T}}}{4\pi r}e^{j(n_{x}+n_{y})\pi}\mathrm{sinc}{\left(\kappa_{x}^{n}\frac{L_{x}}{2}\right)}\mathrm{sinc}{\left(\kappa_{y}^{n}\frac{L_{y}}{2}\right)},$ (11)
where $\kappa_{x}^{n}=\kappa\left(\sin{\theta}\cos{\psi}+\lambda\frac{n_{x}}{L_{x}}\right)$ and $\kappa_{y}^{n}=\kappa\left(\sin{\theta}\sin{\psi}+\lambda\frac{n_{y}}{L_{y}}\right)$ are wavenumbers along the $x$-axis and $y$-axis, respectively.
Proof:
See Appendix -A. ∎
The above proposition provides an explicit expression for Green’s function Fourier transform on the far-field, which we leverage to substitute the integral terms in communication and sensing SINRs.
@@ III-B Representation of Communication and sensing SINRs
$\centering\hat{e}_{k}=j\kappa Z_{0}\sum_{i=1}^{K+M}\sum_{n}^{\infty}x_{i}w_{i,n}f_{2D}^{n}\left(G_{k}\right)+v_{c},~{}k=1,...,K.$ (12)
According to [12, 35], the channel gain $\sum_{n}^{\infty}f_{2D}^{n}\left(G_{k}\right)$ within the band of $[-\kappa,\kappa]$ dominates the radiation power over the Fourier wavenumber domain. Based on Proposition 1, the wavenumber $(\kappa_{x}^{n},\kappa_{y}^{n})$ is determined by the Fourier expansion order $n=(n_{x},n_{y})$. Thus, when the corresponding wavenumbers of the truncation order $N$ satisfy $\kappa_{x}^{n}\geq\kappa$ and $\kappa_{y}^{n}\geq\kappa$, the Fourier series $f_{2D}^{n}\left(G_{k}\right)$ can be approximated up to the $N$-th order, comprising vector
$\centering\boldsymbol{f}_{k}=[f_{2D}^{1}\left(G_{k}\right),...,f_{2D}^{N}\left(G_{k}\right)]^{H},~{}k=1,...,K.$ (13)
@@
$\centering\boldsymbol{w}_{i}=[w_{i,1},...,w_{i,N}]^{H},~{}i=1,...,K+M.$ (14)
$\centering\eta_{c}(\boldsymbol{W};k)=\frac{\boldsymbol{f}_{k}^{H}\boldsymbol{w}_{k}\boldsymbol{w}^{H}_{k}\boldsymbol{f}_{k}}{\boldsymbol{f}_{k}^{H}\boldsymbol{W}\boldsymbol{W}^{H}\boldsymbol{f}_{k}-\boldsymbol{f}_{k}^{H}\boldsymbol{w}_{k}\boldsymbol{w}^{H}_{k}\boldsymbol{f}_{k}+\frac{\sigma_{c}^{2}}{\kappa^{2}Z^{2}_{0}}},~{}k=1,...,K.$ (15)
$\centering\eta_{r}(\boldsymbol{W},\boldsymbol{q}_{l};l)=\frac{\boldsymbol{q}_{l}^{H}\boldsymbol{G}_{l}\boldsymbol{WW}^{H}\boldsymbol{G}^{H}_{l}\boldsymbol{q}_{l}}{\boldsymbol{q}_{l}^{H}\left(\sum_{m=1}^{M}\boldsymbol{G}_{m}\boldsymbol{WW}^{H}\boldsymbol{G}^{H}_{m}-\boldsymbol{G}_{l}\boldsymbol{WW}^{H}\boldsymbol{G}^{H}_{l}\right)\boldsymbol{q}_{l}+\frac{\sigma_{r}^{2}}{\kappa^{2}Z^{2}_{0}}},~{}l=1,...,M.$ (16)
Then, the beamforming matrix $\boldsymbol{W}$ of the whole pattern is expressed by
$\centering\boldsymbol{W}=[\boldsymbol{W}_{c},\boldsymbol{W}_{r}],$ (17)
where $\boldsymbol{W}_{c}\triangleq[\boldsymbol{w}_{1},...,\boldsymbol{w}_{K}]$ and $\boldsymbol{W}_{r}\triangleq[\boldsymbol{w}_{K+1},...,\boldsymbol{w}_{K+M}]$ are the beamforming matrices for communication data streams and sensing streams, respectively. By combining equations (12), (13), (14), and (17), we approximate the received signal at the $k$-th user as
@@
$+\underbrace{j\kappa Z_{0}\boldsymbol{f}_{k}^{H}\boldsymbol{W}_{r}\boldsymbol{s}}_{\mathrm{Sensing~{}Interference}}+v_{c},~{}k=1,...,K.$
Therefore, the SINR at the $k$-th user is expressed by (15).
Similar to the transmit end, the receive pattern in the continuous space domain is converted into the discrete Fourier domain by $\tau_{l}(p^{\prime})=\sum_{n}^{\infty}q_{l,n}\Phi_{n}(p)$, where $q_{l,n}$ is the coefficient of the $l$-th receive pattern projected to $\Phi_{n}(p)$. We confine the Fourier series $g_{2D}^{n}\left(G^{*}_{m}\right)$ over surface $\mathcal{P}_{R}$ to the $N$-th order and comprise vector $\boldsymbol{g}_{m}^{*}=[g_{2D}^{1}\left(G_{m}^{*}\right),...,g_{2D}^{N}\left(G_{m}^{*}\right)]^{H}$. The coefficients of the $l$-th pattern are collected into a vector $\boldsymbol{q}_{l}=[q_{l,1},...,q_{l,N}]^{H}$. Combining (11) with (5), the received echo w.r.t. the $l$-th target is approximated as
@@
$+\underbrace{\boldsymbol{q}_{l}^{H}\boldsymbol{v}_{r}}_{\mathrm{Beamformer~{}Noise}},~{}l=1,...,M,$
where $\boldsymbol{G}_{m}=\boldsymbol{g}_{m}\boldsymbol{g}_{m}^{H}$, and $\boldsymbol{v}_{r}\sim\mathcal{CN}\left(0,\sigma_{r}^{2}\boldsymbol{I}_{N}\right)$ is the Fourier transform of ${v}_{r}$ over surface $D$.
With unit-power beamformer $\boldsymbol{q}_{l}$ at $\mathcal{P}_{R}$, the corresponding sensing SINR w.r.t. the $l$-th target is then expressed as (16).
III-C Problem Formulation in Discrete Domain
From (15) and (16), both the sensing and communication SINRs are converted into the discrete domain and determined by transmit beamformers in $\boldsymbol{W}$ and receive beamformers collected in $\boldsymbol{Q}=[\boldsymbol{q}_{1},...,\boldsymbol{q}_{M}]$. The original problem in (7) is converted from the continuous domain into the discrete domain. Based on the transformation in Section III-A and truncation in Section III-B, the resulting approximated problem w.r.t. the joint transmit-receive beamforming can be formulated as

### v2

@@ II-B Communications and Radar Performances
$+\sum_{i=1,i\neq k}^{K+M}j\kappa Z_{0}\int_{D}G_{k}(p)\rho_{i}(p)x_{i}\mathrm{d}p+v_{c},~k=1,.,K,$
where $v_{c}$ is the additive white Gaussian noise (AWGN) with variance $\sigma_{c}^{2}$, $\kappa$ is the wavenumber, and $Z_{0}$ is the wave impedance in the free space. The communications SINR of the $k$-th user, for k=1,,Kk=1,...,K, is given by
$\centering\eta_{c}(\boldsymbol{\rho}(p);k)=\frac{\mathbf{E}\left(\left|j\kappa Z_{0}\int_{D}G_{k}(p)\rho_{k}(p)c_{k}\mathrm{d}p\right|^{2}\right)}{\mathbf{E}\left(\left|\sum_{i=1,i\neq k}^{K+M}j\kappa Z_{0}\int_{D}G_{k}(p)\rho_{i}(p)x_{i}\mathrm{d}p\right|^{2}\right)+\sigma_{c}^{2}},$ (3)
In our considered HISAC system, both communication data streams and radar streams are utilized to support multi-target sensing since they are known by the $\mathcal{P}_{R}$ from prior sensing steps. In particular, we consider sensing SINR to design a focused HIS beampattern toward regions of interest identified in a previous sensing stage, thereby enabling more refined sensing tasks. Following the standard full-duplex ISAC assumption [2, 3, 1, 4, 23], self-interference (SI) after cancellation is assumed to be below the noise floor and is therefore omitted [37, 38]. Considering both transmit and receive patterns, the sensing SINR is formulated in the following.
Let $G_{m}(p)\triangleq G\left(p^{\prime}_{m},p\right)$ and $G_{m}^{*}(p)\triangleq G\left(p,p^{\prime}_{m}\right)$ denote the Green’s function from source $p$ to $m$-th targets and from the $m$-th targets to source $p$, respectively. The $m$-th target’s echo at $p^{\prime}$ is expressed as
$\centering y_{m}(p^{\prime})=\alpha_{m}G_{m}^{*}(p)e_{m},~m=1,.,M,~p^{\prime}\in\mathcal{P}_{R},$ (4)
@@
$\centering y_{l}=$ $\sum_{m=1}^{M}\int_{D}\tau_{l}\left(p^{\prime}\right)y_{m}\left(p^{\prime}\right)\mathrm{d}p^{\prime}+\int_{D}\tau_{l}(p^{\prime})v_{r}\left(p^{\prime}\right)\mathrm{d}p^{\prime},$ (5)
where ${v}_{r}(p^{\prime}),~p^{\prime}\in D$ is the AWGN at $\mathcal{P}_{R}$ with variance $\sigma_{r}^{2}$. Therefore, for l=1,,Ml=1,...,M, the sensing SINR is given by
$\centering\eta_{r}(\boldsymbol{\rho},\boldsymbol{\tau};l)=\frac{\mathbf{E}\left(\left|e_{l}\int_{D}\tau_{l}\left(p^{\prime}\right)G_{l}^{*}(p)\mathrm{d}p^{\prime}\right|^{2}\right)}{\mathbf{E}\left(\left|\sum_{m=1,m\neq l}^{M}e_{m}\int_{D}\tau_{l}\left(p^{\prime}\right)G_{m}^{*}(p)\mathrm{d}p^{\prime}\right|^{2}\right)+\sigma_{r}^{2}}.$ (6)
@@ II-C Problem Formulation in Continuous Domain
$\ \int_{D}\|\boldsymbol{\rho}\|_{2}^{2}\mathrm{d}p\leq P_{T},$ (7c)
$\ \int_{D}|{\tau}_{l}|^{2}\mathrm{d}p^{\prime}=1,~l=1,...,M.$ (7d)
The primary goal of (7) is to balance the sensing SINR among radar targets, and (7b) is the communication constraint for multiple users, where $\Gamma_{c}$ represents the minimal required communication SINR. Concerning power constraints, (7c) and (7d) represent the power budget for the transmit pattern at $\mathcal{P}_{S}$ and the unit-power constraint for the receive pattern at $\mathcal{P}_{R}$, respectively. Generally, the receive beamformer’s power in the direction of interest is maintained at $1$ to ensure a distortionless response for the signal coming from that specific direction, where receive beamforming vectors satisfy $|{\tau}_{l}|^{2}\mathrm{d}p^{\prime}=1$.
Due to the presence of infinite-dimensional integrals in both the objective and constraints of problem (7), classical beamforming techniques designed for discrete-array-based ISAC systems are not directly applicable. Moreover, as illustrated in Fig. 2, individually designing each current cell on the HIS surface is impractical from a hardware perspective, motivating the need for a more feasible alternative. To address the challenges posed by the infinite-dimensional functions in problem (7), we develop a continuous-discrete transformation inspired by the Fourier transform and propose a tractable hardware implementation of the HIS, as detailed in the following section.
@@ III Continuous-Discrete Transformation
In this section, we formulate a comprehensive continuous-discrete transformation for $\boldsymbol{\rho}$, which is then applied to derive the expressions for communication and sensing SINRs. Based on this transformation, we approximate the continuous pattern design in (7) into a joint transmit-receive beamforming problem in the discrete domain.
@@ III-A Comprehensive Continuous-Discrete Transformation
To overcome the integral functions in the pattern design, we provide a closed-form expression for $f_{2D}^{n}\left(G_{p^{\prime}}\right)$ in the following.
The expression of the scalar Green’s function in the far-field is given by
$\centering G(\textbf{r},\textbf{s})=\frac{e^{j\kappa\|\textbf{r}-\textbf{s}\|}}{4\pi\|\textbf{r}-\textbf{s}\|},$ (10)
$\centering\|\textbf{r}-\textbf{s}\|\approx r-\sin{\theta}\left(s_{x}\cos{\psi}+s_{y}\sin{\psi}\right).$ (11)
By substituting (11) into (10), the Green’s function on the far field is approximated as
$\centering G_{p^{\prime}}=\frac{e^{j\kappa r}}{4\pi r}e^{-j\kappa s_{x}\sin{\theta}\cos{\psi}}e^{-j\kappa s_{y}\sin{\theta}\sin{\psi}}.$ (12)
Combining (12) and (9), the Green’s function Fourier transform $f_{2D}^{n}\left(G_{p^{\prime}}\right)$, is given by
$\centering f_{2D}^{n}\left(G_{p^{\prime}}\right)=\int_{D}\frac{e^{j\kappa r}}{4\pi r}e^{-j\kappa s_{x}\sin{\theta}\cos{\psi}}e^{-j\kappa s_{y}\sin{\theta}\sin{\psi}}\Psi_{n}(p)\mathrm{d}p,$ (13)
$\centering\Psi_{n}(p)=\frac{1}{\sqrt{A_{T}}}e^{-j2\pi\left(\frac{n_{x}}{L_{x}}\left(s_{x}-\frac{L_{x}}{2}\right)\right)}e^{-j2\pi\left(\frac{n_{y}}{L_{y}}\left(s_{y}-\frac{L_{y}}{2}\right)\right)}.$ (14)
Let $\kappa_{x}^{n}=\kappa\left(\sin{\theta}\cos{\psi}+\lambda\frac{n_{x}}{L_{x}}\right)$ and $\kappa_{y}^{n}=\kappa\left(\sin{\theta}\sin{\psi}+\lambda\frac{n_{y}}{L_{y}}\right)$ denote the wavenumbers along the $x$-axis and $y$-axis, respectively. Combining (13) with (14), the closed-form Green’s function Fourier transform for arbitrary position $(r,\theta,\psi)$ can be calculated by (15).
$\centering f_{2D}^{n}$ $\left(G_{p^{\prime}}\right)=\frac{e^{j\kappa r}}{4\pi r\sqrt{A_{T}}}e^{j(n_{x}+n_{y})\pi}$ (15)
$\int_{L_{x}}e^{-j\kappa_{x}^{n}s_{x}}\mathrm{d}s_{x}\int_{L_{y}}e^{-j\kappa_{y}^{n}s_{y}}\mathrm{d}s_{y}$
$=$ $\frac{e^{j\kappa r}e^{j(n_{x}+n_{y})\pi}}{4\pi r\sqrt{A_{T}}}\left[\frac{e^{-j\kappa_{x}^{n}s_{x}}}{-j\kappa_{x}^{n}}\right]_{-\frac{L_{x}}{2}}^{\frac{L_{x}}{2}}\left[\frac{e^{-j\kappa_{y}^{n}s_{y}}}{-j\kappa_{y}^{n}}\right]_{-\frac{L_{y}}{2}}^{\frac{L_{y}}{2}}$
It’s obvious that $f_{2D}^{n}\left(G_{p^{\prime}}\right)$ is determined by the $\kappa_{x}^{n}$ and $\kappa_{y}^{n}$. Taking the $x$-axis as example, when the $x$-axis wavenumber satisfies
$\centering\kappa\left(\sin{\theta}\cos{\psi}+\lambda\frac{n_{x}}{L_{x}}\right)\cdot\frac{L_{x}}{2}\geq n\pi,$ (16)
the corresponding $\mathrm{sinc}$ term approaches its first zero, indicating that higher-order terms contribute negligibly to the total response and can be safely truncated. Consequently, the truncation order $(n_{x},n_{y})$ should satisfy $\{|n_{x}|\geq\frac{L_{x}}{\lambda};|n_{y}|\geq\frac{L_{y}}{\lambda}\}$, and the minimum total truncation order is bounded by $N\geq\left(2\frac{L_{x}}{\lambda}+1\right)\left(2\frac{L_{y}}{\lambda}+1\right)$. It is worth clarifying that the spatial DoFs of the aperture are fundamentally determined by its electrical size [39] and do not depend on the truncation order itself. In particular, the maximum number of independent spatial DoFs supported by the continuous-aperture is bounded by $\frac{\pi A_{T}}{\lambda^{2}}$. The condition in (16) specifies a sufficient truncation order such that the retained Fourier modes fully capture the intrinsic DoFs supported by the continuous-aperture. When this condition is satisfied, higher-order terms contribute negligibly, and increasing the truncation order beyond this point does not yield additional DoFs.
@@ III-B Representation of Communication and sensing SINRs
$\centering\hat{e}_{k}=j\kappa Z_{0}\sum_{i=1}^{K+M}\sum_{n}^{\infty}x_{i}w_{i,n}f_{2D}^{n}\left(G_{k}\right)+v_{c},~k=1,.,K.$ (17)
According to [9, 34], the channel gain $\sum_{n}^{\infty}f_{2D}^{n}\left(G_{k}\right)$ within the band of $\kappa,\kappa]$ dominates the radiation power over the Fourier wavenumber domain. According to (16), the Fourier series $f_{2D}^{n}\left(G_{k}\right)$ can be approximated up to the $N$-th order, comprising vector
$\centering\boldsymbol{f}_{k}=[f_{2D}^{1}\left(G_{k}\right),.,f_{2D}^{N}\left(G_{k}\right)]^{H},~k=1,.,K.$ (18)
@@
$\centering\boldsymbol{w}_{i}=[w_{i,1},.,w_{i,N}]^{H},~i=1,.,K+M.$ (19)
Then, the beamforming matrix $\boldsymbol{W}$ of the whole pattern is expressed by $\boldsymbol{W}=[\boldsymbol{W}_{c},\boldsymbol{W}_{r}]$, where $\boldsymbol{W}_{c}\triangleq[\boldsymbol{w}_{1},...,\boldsymbol{w}_{K}]$ and $\boldsymbol{W}_{r}\triangleq[\boldsymbol{w}_{K+1},...,\boldsymbol{w}_{K+M}]$ are the beamforming matrices for communication data streams and sensing streams, respectively. By combining equations (17), (18), and (19), we approximate the received signal at the $k$-th user as
@@
$+\underbrace{j\kappa Z_{0}\boldsymbol{f}_{k}^{H}\boldsymbol{W}_{r}\boldsymbol{s}}_{\mathrm{Sensing~Interference}}+v_{c},~k=1,.,K.$
Therefore, combining (3) with (20), the SINR at the $k$-th user is expressed by (21).
$\centering\eta_{c}(\boldsymbol{W};k)=\frac{\boldsymbol{f}_{k}^{H}\boldsymbol{w}_{k}\boldsymbol{w}^{H}_{k}\boldsymbol{f}_{k}}{\boldsymbol{f}_{k}^{H}\boldsymbol{W}\boldsymbol{W}^{H}\boldsymbol{f}_{k}-\boldsymbol{f}_{k}^{H}\boldsymbol{w}_{k}\boldsymbol{w}^{H}_{k}\boldsymbol{f}_{k}+\frac{\sigma_{c}^{2}}{\kappa^{2}Z^{2}_{0}}},~k=1,.,K.$ (21)
$\centering\eta_{r}(\boldsymbol{W},\boldsymbol{q}_{l};l)=\frac{\boldsymbol{q}_{l}^{H}\boldsymbol{G}_{l}\boldsymbol{WW}^{H}\boldsymbol{G}^{H}_{l}\boldsymbol{q}_{l}}{\boldsymbol{q}_{l}^{H}\left(\sum_{m=1}^{M}\boldsymbol{G}_{m}\boldsymbol{WW}^{H}\boldsymbol{G}^{H}_{m}-\boldsymbol{G}_{l}\boldsymbol{WW}^{H}\boldsymbol{G}^{H}_{l}\right)\boldsymbol{q}_{l}+\frac{\sigma_{r}^{2}}{\kappa^{2}Z^{2}_{0}}},~l=1,.,M.$ (22)
Similar to the transmit end, the receive pattern in the continuous space domain is converted into the discrete Fourier domain by $\tau_{l}(p^{\prime})=\sum_{n}^{\infty}q_{l,n}\Phi_{n}(p^{\prime})$, where $q_{l,n}$ is the coefficient of the $l$-th receive pattern projected to $\Phi_{n}(p^{\prime})$. We confine the Fourier series $g_{2D}^{n}\left(G^{*}_{m}\right)$ over surface $\mathcal{P}_{R}$ to the $N$-th order and comprise vector $\boldsymbol{g}_{m}^{*}=[g_{2D}^{1}\left(G_{m}^{*}\right),...,g_{2D}^{N}\left(G_{m}^{*}\right)]^{H}$. The coefficients of the $l$-th pattern are collected into a vector $\boldsymbol{q}_{l}=[q_{l,1},...,q_{l,N}]^{H}$. Combining (15) with (5), the received echo w.r.t. the $l$-th target is approximated as
@@
$+\underbrace{\boldsymbol{q}_{l}^{H}\boldsymbol{v}_{r}}_{\mathrm{Beamformer~Noise}},~l=1,.,M,$
where $\boldsymbol{G}_{m}=\boldsymbol{g}_{m}\boldsymbol{g}_{m}^{H}$, and $\boldsymbol{v}_{r}$ is the Fourier transform of ${v}_{r}$ over surface $D$ via the truncated oreder $N$.
Remark 1
The discrete-domain receive noise vector $\boldsymbol{v}_{r}$ obtained via the truncated Fourier representation follows an i.i.d. circularly symmetric complex Gaussian distribution with covariance $\boldsymbol{v}_{r}\sim\mathcal{CN}\left(0,\sigma_{r}^{2}\boldsymbol{I}_{N}\right)$.
Proof:
See Appendix. -B
Accordingly, the receive beamformer $\boldsymbol{q}_{l}$ of the HISAC system ensures that the signal of interest is received without amplification and attenuation, allowing for accurate reception and signal processing. With unit-power beamformer $\boldsymbol{q}_{l}$ at $\mathcal{P}_{R}$, the corresponding sensing SINR w.r.t. the $l$-th target in (6) is then expressed as (22).
Figure 3: The fully-digital HIS architecture.
III-C Antenna Architecture
According to the continuous-discrete transformation, the HIS beam pattern can be reconstructed by summing $N$ beams with specific wavenumbers. Consequently, the continuous beam pattern of the HIS can be effectively approximated using a discrete set of $N$ beams. To achieve the continuous-discrete transformation of the HIS and optimize its pattern, we consider the fully-digital architecture, as shown in Fig. 3. In the following subsections, we briefly introduce the antenna architecture and discuss its corresponding beamformer models.
III-C1 Fully-Digital Architecture
In the fully-digital antenna architecture, each beam’s feed is connected to a dedicated RF chain, as shown in Fig. 3. This configuration provides the HISAC system with maximum flexibility in signal processing, allowing independent control over the patterns of each beam. Within this framework, the transmit HIS beamforming design corresponds directly to the Fourier transform coefficient matrix $\boldsymbol{W}$ in (21) and (22). Similarly, the receive HIS beamforming is implemented through simultaneous reception, which aligns with the coefficient vectors $\boldsymbol{q}_{l}$ in (22). Consequently, leveraging the principle of continuous-discrete transformation, HIS’s beam can be turned by designing the $\boldsymbol{W}$ matrix within the fully-digital architecture.
To implement the continuous HIS in hardware, we construct the surface using $N=N_{x}\times N_{y}$ square-shaped radiators, densely tiled to form the full aperture. Here, $N_{x}$ and $N_{y}$ denote the number of truncated Fourier orders along the $x$- and $y$-axes, respectively. Each radiator occupied an area of $A_{T}/(N_{x}-1)(N_{y}-1)$ serves as a sub-element that supports programmable radiation or reception characteristics. On the transmit side, the radiation function of each unit is operated by the basis function $f^{n}_{2D}(\cdot)$, while on the receive side, the directional response is represented by $g^{n}_{2D}(\cdot)$. The overall EM behavior of the HIS is then synthesized through the weighted superposition of these $N$ radiators.
Unlike discrete arrays that use dipole-like isotropic radiators, HIS employs a continuous surface composed of densely packed rectangular patches to excite surface currents [13, 16], enabling enhanced beamforming control. Note that the number of RF chains in our fully-digital architecture is set to be equal to the truncation order $N$, ensuring one-to-one mapping between digital control and spatial basis functions. In future work, the proposed design may be extended to hybrid architectures with fewer RF chains than $N$, where spatial multiplexing can be achieved by phase shifters.
III-D Problem Formulation in Discrete Domain
Based on the above antenna architecture, we consider beamforming matrices $\boldsymbol{W}$ and $\boldsymbol{Q}$ as optimization variables. From (21) and (22), both the sensing and communication SINRs are converted into the discrete domain and determined by transmit beamformers in $\boldsymbol{W}$ and receive beamformers collected in $\boldsymbol{Q}=[\boldsymbol{q}_{1},...,\boldsymbol{q}_{M}]$. The original problem in (7) is converted from the continuous domain into the discrete domain. Based on the transformation in Section III-A and truncation in Section III-B, the resulting approximated problem w.r.t. the joint transmit-receive beamforming can be formulated as